# Multi-Pair Temporal Sentence Grounding via Multi-Thread Knowledge Transfer Network

### v1

@@ Introduction
Temporal sentence grounding (TSG) (Gao et al. 2017; Fang et al. 2022, 2023b, 2023c, 2024c, 2023a, 2024d; Li et al. 2023, 2024; Yu et al. 2024c; Ning et al. 2024, 2023a, 2023b) is an important yet challenging multi-modal task, which has received increasing attention in recent years due to its wide potential applications, such as video understanding (Liu et al. 2024c, 2023e, d, b; Wang et al. 2024c; Hu et al. 2023; Fei et al. 2024d, c, b, a; Wu et al. 2024; Wang et al. 2024b, 2023, 2022a; Zhang et al. 2024a, b; Yu et al. 2022) and human-computer interaction (Liu et al. 2024e, f, 2023g, 2023c, 2023d; Tang et al. 2022b, a, 2023a, 2024a; Feng et al. 2023b, a; Zhao et al. 2024a, b; Wang et al. 2024a; Yu et al. 2024b, 2023). By complex multi-modal interactions and complicated context information, TSG targets the challenging problem of locating a variety of sentence queries about a video, which requires the designed models to understand both natural language and long video, including reasoning about activities, objects, sequence of events, and interactions within the video (Hu et al. 2022; Wei et al. 2023; Zhao et al. 2021, 2017, 2022, 2018a, 2018b; Jia et al. 2019, 2022b, 2022c, 2022a, 2024a, 2021, 2024d, 2020, 2023, 2024b, 2024e, 2024c; Gao et al. 2021, 2022, 2024a, 2024b, 2024c; Fang et al. 2024a). As shown in Figure 1(a), given an untrimmed video and a sentence query, TSG aims to determine the segment boundaries that contain the query-relevant activity (Qu et al. 2020; Dong et al. 2022b; Sun et al. 2024; Qu et al. 2024a; Zheng et al. 2023; Qu et al. 2024b).
@@
Most previous TSG works (Xiu et al. 2024; Liu et al. 2022, 2020, 2021; Ji et al. 2023c, b, 2024d, 2024b, 2024a, a) refer to a fully-supervised setting, where each frame is manually labeled as query-relevant or not. To avoid using such labour-intensive frame-level annotations, some recent works (Fang et al. 2025, 2020; Fang and Hu 2020; Fang et al. 2021b, a; Fang, Easwaran, and Genest 2024; Fang et al. 2024b; Liu et al. 2024i, 2023f, g, h; Zhang et al. 2024c; Yu et al. 2024a; Zhang et al. 2024d, 2023d, 2023c, 2023b, 2022; Guo et al. 2024; Dong et al. 2022a) explore a weakly-supervised setting with only the video-query correspondence to alleviate the reliance to a certain extent. Despite the remarkable performance, fully- and weakly-supervised methods only treat each video-query pair independently and ignore the semantic relationship between different video-query pairs, as shown in Figure 1(b). Since a video often corresponds to multiple queries, their treatment will repeatedly extract video features and repetitively conduct complex multi-modal calculations, resulting in weak efficiency. Besides, ignoring the semantic relationship between different video-query pairs might miss important spatio-temporal semantic consistency (e.g., common noun/appearance “woman” and temporal relationship “continues to” in Figure 2), which limits their effectiveness. Thus, an effective and efficient model is expected to explore the latent semantics relationship between different video-query pairs.
@@
Fully-supervised TSG. Temporal sentence grounding (TSG) (Anne Hendricks et al. 2017; Gao et al. 2017; Ju et al. 2024, 2020, 2023; Liang et al. 2024a, b, 2023) aims at locating the most relevant segments from long videos corresponding to the given sentence descriptions. Traditional TSG methods (Gao et al. 2017; Xiong et al. 2024; Rao et al. 2021b, a; Jiang et al. 2023; Zhang et al. 2023a; Wang et al. 2025b, a; Tang et al. 2024d, 2022c, 2021, 2022d, 2023b, b, c) typically utilize a propose-and-rank approach to make predictions based on interacted multi-modal features. Some proposal-free methods (Zhang et al. 2020a; Lin et al. 2024b, a, c, 2025; Wang et al. 2021a, b, 2020) are proposed to directly predict the temporal locations of the target segment without generating proposals.
@@
Although previous TSG works (Anne Hendricks et al. 2017; Gao et al. 2017) try to fully understand textual information and visual information, they often ignore the semantic relationship between different sentences. The semantic relationship includes 1) the temporal information between different segments in the same video, and 2) the contextual information among different sentences. To sufficiently mine these query-to-query relationships, we extract query-level features $F_{q}=\{q_{e}^{j}\}_{j=1}^{M_{q}}$ from the multiple queries rather than learnable embeddings in previous works. Thus, we aim to model the query-level contexts, and explore the temporally and contextually related queries of each query. For example, in Figure 2, “The woman starts doing some ballet moves.” and “She continues to dance to ballet moves as the little girl walks away.” share the same semantics (“woman” and “ballet moves”), and contains the temporal relationship (“start” and “continues to”). The shared semantics and temporal relationship will assist the grounding task of each sentence.
@@
Compared methods. 1) Fully-supervised (FS): DRN (Zeng et al. 2020), 2D-TAN (Zhang et al. 2020b), MIGCN (Zhang et al. 2021), MMN (Wang et al. 2022c), VDI (Luo et al. 2023), MRTNet (Ji et al. 2024c), MESM (Liu et al. 2024j). 2) Weakly-supervised (WS): LCNet (Yang et al. 2021), VCA (Wang, Chen, and Jiang 2021), ICVC (Chen et al. 2022), WSTAN (Wang et al. 2022b), CNM (Zheng et al. 2022a), CPL (Zheng et al. 2022b), MMDist (Bao et al. 2024).

### v2

@@ Introduction
Temporal sentence grounding (TSG) (Gao et al. 2017; Fang et al. 2022, 2023b, 2023c, 2024b, 2023a, 2024c; Li et al. 2023, 2024; Yu et al. 2024c; Ning et al. 2024, 2023a, 2023b; Wang et al. 2025a) is an important yet challenging multi-modal task, which has received increasing attention in recent years due to its wide potential applications, such as video understanding (Liu et al. 2024c, 2023e, d, b; Wang et al. 2024c; Hu et al. 2023a; Fei et al. 2024d, c, b, a; Wu et al. 2024; Wang et al. 2024b, 2023, 2022a; Zhang et al. 2024a, b; Yu et al. 2022, 2025) and human-computer interaction (Liu et al. 2024f, g, 2023g, 2023c, 2023d; Tang et al. 2022b, a, 2023a, 2024a; Feng et al. 2023b, a; Zhao et al. 2024a, b; Wang et al. 2024a; Yu et al. 2024b, 2023). By complex multi-modal interactions and complicated context information, TSG targets the challenging problem of locating a variety of sentence queries about a video, which requires the designed models to understand both natural language and long video, including reasoning about activities, objects, sequence of events, and interactions within the video (Hu et al. 2022; Wei et al. 2023; Zhao et al. 2021, 2017, 2022, 2018a, 2018b; Jia et al. 2019, 2022b, 2022c, 2022a, 2024a, 2021, 2024d, 2020, 2023, 2024b, 2024e, 2024c; Gao et al. 2021, 2022, 2024a, 2024b, 2024c; Fang et al. 2025a; Hu et al. 2020b, 2021b, 2023b, 2023b; Hu, Liu, and Wu 2021; Hu et al. 2020c, 2024d, a, 2024d, 2023d, 2024e, 2024c, 2021a, 2024b, 2024e; Hu and Su 2024; Hu et al. 2024e, 2023f, 2023e, 2023c; Li et al. 2022, 2025). As shown in Figure 1(a), given an untrimmed video and a sentence query, TSG aims to determine the segment boundaries that contain the query-relevant activity (Qu et al. 2020; Dong et al. 2022b; Sun et al. 2024; Qu et al. 2024a; Zheng et al. 2023; Qu et al. 2024b).
@@
Most previous TSG works (Xiu et al. 2024; Liu et al. 2022, 2020, 2021; Ji et al. 2023c, b, 2024d, 2024b, 2024a, a) refer to a fully-supervised setting, where each frame is manually labeled as query-relevant or not. To avoid using such labour-intensive frame-level annotations, some recent works (Fang et al. 2025b, 2020; Fang and Hu 2020; Fang et al. 2021b, a; Fang, Easwaran, and Genest 2024; Fang et al. 2024a; Liu et al. 2024j, 2023f, h, i; Zhang et al. 2024c; Yu et al. 2024a; Zhang et al. 2024d, 2023d, 2023c, 2023b, 2022; Guo et al. 2024; Dong et al. 2022a) explore a weakly-supervised setting with only the video-query correspondence to alleviate the reliance to a certain extent. Despite the remarkable performance, fully- and weakly-supervised methods only treat each video-query pair independently and ignore the semantic relationship between different video-query pairs, as shown in Figure 1(b). Since a video often corresponds to multiple queries, their treatment will repeatedly extract video features and repetitively conduct complex multi-modal calculations, resulting in weak efficiency. Besides, ignoring the semantic relationship between different video-query pairs might miss important spatio-temporal semantic consistency (e.g., common noun/appearance “woman” and temporal relationship “continues to” in Figure 2), which limits their effectiveness. Thus, an effective and efficient model is expected to explore the latent semantics relationship between different video-query pairs.
@@
Fully-supervised TSG. Temporal sentence grounding (TSG) (Anne Hendricks et al. 2017; Gao et al. 2017; Ju et al. 2024, 2020, 2023; Liang et al. 2024a, b, 2023) aims at locating the most relevant segments from long videos corresponding to the given sentence descriptions. Traditional TSG methods (Gao et al. 2017; Xiong et al. 2024; Rao et al. 2021b, a; Jiang et al. 2023; Zhang et al. 2023a; Wang et al. 2025c, b; Tang et al. 2024d, 2022c, 2021, 2022d, 2023b, b, c) typically utilize a propose-and-rank approach to make predictions based on interacted multi-modal features. Some proposal-free methods (Zhang et al. 2020a; Lin et al. 2024b, a, c, 2025; Wang et al. 2021a, b, 2020) are proposed to directly predict the temporal locations of the target segment without generating proposals.
@@
Although previous TSG works (Anne Hendricks et al. 2017; Gao et al. 2017) try to fully understand textual information and visual information (Zhao et al. 2024a; Xiong et al. 2024; Cai et al. 2025; Liu et al. 2024e; Tang et al. 2025, 2024d; Lei et al. 2025; Yang et al. 2025; Liu et al. 2024g; Zhang et al. 2025a, b; Hu et al. 2024a, 2025), they often ignore the semantic relationship between different sentences. The semantic relationship includes 1) the temporal information between different segments in the same video, and 2) the contextual information among different sentences. To sufficiently mine these query-to-query relationships, we extract query-level features $F_{q}=\{q_{e}^{j}\}_{j=1}^{M_{q}}$ from the multiple queries rather than learnable embeddings in previous works. Thus, we aim to model the query-level contexts, and explore the temporally and contextually related queries of each query. For example, in Figure 2, “The woman starts doing some ballet moves.” and “She continues to dance to ballet moves as the little girl walks away.” share the same semantics (“woman” and “ballet moves”), and contains the temporal relationship (“start” and “continues to”). The shared semantics and temporal relationship will assist the grounding task of each sentence.
@@
Compared methods. 1) Fully-supervised (FS): DRN (Zeng et al. 2020), 2D-TAN (Zhang et al. 2020b), MIGCN (Zhang et al. 2021), MMN (Wang et al. 2022c), VDI (Luo et al. 2023), MRTNet (Ji et al. 2024c), MESM (Liu et al. 2024e). 2) Weakly-supervised (WS): LCNet (Yang et al. 2021), VCA (Wang, Chen, and Jiang 2021), ICVC (Chen et al. 2022), WSTAN (Wang et al. 2022b), CNM (Zheng et al. 2022a), CPL (Zheng et al. 2022b), MMDist (Bao et al. 2024).